# ALPy Cogenesis

### v1

@@ I Introduction
The origin of dark matter (DM) and baryon asymmetry in the Universe (BAU) have been longstanding puzzles in particle physics and cosmology Zyla:2020zbs ; Planck:2018vyg . The failure of the standard model (SM) of particle physics to explain these observed phenomena has led to several beyond standard model (BSM) proposals in the literature. Among them, the weakly interacting massive particle (WIMP) paradigm of DM Kolb:1990vq ; Jungman:1995df and baryogenesis/leptogenesis Weinberg:1979bt ; Kolb:1979qa ; Fukugita:1986hr are the most popular ones. While the minimal WIMP scenarios are pushed to a tight corner by null results at direct search experiments LZ:2022lsv , generic baryogenesis models operate at a high scale out of reach from direct experimental probe. Minimal vanilla leptogenesis scenarios also has a lower bound on the scale $\geq 10^{9}$ GeV, known as the Davidson-Ibarra (DI) bound Davidson:2002qv . While there exists low scale realisation of leptogenesis with resonantly enhanced CP asymmetry Pilaftsis:2003gt , some well-motivated DM paradigms beyond the simplest WIMP paradigm have also been explored in the literature. Axion or axion like particle (ALP) provides such an alternative paradigm for DM, a recent review of which can be found in OHare:2024nmr .
@@
Axion Wilczek:1977pj ; Weinberg:1977ma is the pseudo Nambu-Goldstone boson (pNGB) associated with the Peccei-Quinn (PQ) symmetry $U(1)_{\rm PQ}$ Peccei:1977hh ; Peccei:1977ur ; Wilczek:1977pj ; Weinberg:1977ma introduced to solve the strong CP problem, another longstanding puzzle in particle physics. In addition to solving the strong CP problem, axion with sub-eV mass can constitute the entire cold dark matter in the Universe Preskill:1982cy ; Abbott:1982af ; Dine:1982ah . Axions can also have interesting detection prospects due to their couplings to the SM particles, suppressed by the PQ symmetry breaking scale $f$. For QCD axion models with sub-eV axion mass, there exists a stringent astrophysical bound $f\gtrsim 10^{8}-10^{9}$ GeV Raffelt:2006cw ; Caputo:2024oqc . On the other hand, ALP parameter space is less constrained as their coupling and mass can be varied independently.
@@
Motivated by these, we propose a novel way of connecting leptogenesis and axion by considering a specific interaction involving axion and suppressed by the PQ scale $f$ to be the origin of CP violation. While there are several ways of connecting axion with baryogenesis or leptogenesis Kusenko:2014uta ; Servant:2014bla ; Adshead:2015jza ; Ipek:2018lhm ; Croon:2019ugf ; Co:2019wyp ; Ballesteros:2016xej ; Clarke:2015bea ; Co:2020jtv ; Sopov:2022bog ; Domcke:2022kfs ; Borah:2024qyo ; Mishra:2025twb , we for the first time, propose a decay of the type $\psi_{1}\rightarrow\psi_{2}a$ (with $\psi_{1,2}$ being chiral fermions and $a$ being the axion) to be the origin of CP asymmetry. As a first working example of this mechanism, dubbed as ALPy cogenesis, we consider $\psi_{1}$ to be a heavy right-handed neutrino (RHN) $N_{R}$ while $\psi_{2}$ to be a light chiral fermion $\nu_{R}$, both of which are singlets under the SM gauge symmetry [^1] In another recent work Cataldi:2024bcs , a non-thermal leptogenesis scenario was studied where a heavy ALP decays into RHN in the early Universe. On the other hand, Datta:2024xhg considered spontaneous leptogenesis with sub-GeV ALP.. Non-zero CP asymmetry is generated after taking one-loop correction to this decay if at least two copies of $N_{R}$ are considered. The asymmetry generated in $\nu_{R}$ is then transferred to left-handed leptons via sizeable Yukawa interactions with a neutrinophilic Higgs doublet Heeck:2013vha ; Borah:2022qln . Due to the lepton number violation induced by heavy RHN $N_{R}$, a Majorana mass of $\nu_{R}$ is also generated at one-loop level. This radiative correction splits the three Dirac neutrinos at tree level into three active and three sterile Majorana or pseudo-Dirac neutrinos. Keeping the Majorana mass of $\nu_{R}$ small to control the washouts of asymmetry created from $N_{R}$ decay also constrains the scale of leptogenesis and other parameters of the model. As a second working example, we consider a setup without lepton number violation which ensures light neutrinos to be of Dirac type. This requires the presence of another Dirac fermion $\chi$ such that net CP asymmetries in $\nu_{R}$ and $\chi$ can be zero due to lepton number conservation. Keeping the $\nu_{R}$ and $\chi$ sectors out of equilibrium allow a net lepton asymmetry to survive till the sphaleron decoupling epoch. The asymmetric $\chi$, if sufficiently long lived, can contribute to DM along with axions. In both the examples, the scale of leptogenesis as well as the PQ scale are constrained from the requirement of generating the observed BAU and DM. The PQ scale, on the other hand, also controls the detection prospects of axion or ALP over a wide range of experiments. While a thermal leptogenesis leads to thermalised axions, the Dirac ALPy cogenesis also has thermalised Dirac neutrinos both of which lead to enhanced dark radiation within reach of current and future cosmic microwave background (CMB) experiments.
@@
This paper is organised as follows. In section II, we discuss the minimal framwork with Majorana neutrinos. In section III, we discuss the Dirac ALPy cogenesis scenario followed by the detection prospects of both the scenarios in section IV. We finally conclude in section V.
@@ II Minimal Framework
In the minimal realisation of the ALPy cogenesis idea, the SM lepton content is extended by two different types of singlet chiral fermions $N_{R},\nu_{R}$. While $\nu_{R}$ combines with the left chiral neutrinos $\nu_{L}$ to form Dirac neutrinos at tree level via a neutrinophilic Higgs doublet $H_{2}$, the heavy right-handed neutrino $N_{R}$ has an effective axion-like particle (ALP) coupling with $\nu_{R}$. The relevant Lagrangian is given by
$-\mathcal{L}$ $\supset\frac{\partial_{\mu}a}{f}\overline{N_{R_{i}}}(\lambda_{i\alpha}\gamma^{\mu}+\lambda^{\prime}_{i\alpha}\gamma^{\mu}\gamma^{5})\nu_{R_{\alpha}}+\frac{1}{2}M_{i}\overline{N^{c}_{R_{i}}}N_{R_{i}}$
@@
Similar higher dimensional operators with axion interaction being replaced by electromagnetic dipole operator were used for leptogenesis in earlier works Bell:2008fm ; Choudhury:2011gbi ; Borah:2025oqj . Fig. 1 shows the two-body decay $N_{i}\rightarrow\nu_{R_{\alpha}}a$ and possible one-loop corrections responsible for generating the CP asymmetry. The corresponding CP asymmetry can be written as
@@
$\Lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz,\$
and $r_{a}=m_{a}/M_{i},r_{\alpha,\beta}=m_{\nu_{\alpha,\beta}}/M_{i}$, $\widetilde{\lambda}_{i\alpha}=\lambda_{i\alpha}+\lambda^{\prime}_{i\alpha}$. The CP asymmetry parameter, considering self-energy correction, is given by
$\epsilon_{\alpha}=\frac{{\rm Im}[\widetilde{\lambda}_{j\alpha}\widetilde{\lambda}_{j\beta}\widetilde{\lambda}^{*}_{i\beta}\widetilde{\lambda}^{*}_{j\alpha}]}{|\widetilde{\lambda}_{i\beta}|^{2}|\widetilde{\lambda}_{i\alpha}|^{2}}\Gamma_{0}(N_{i}\rightarrow\nu_{R\beta}a)\frac{r_{ij}}{1-r^{2}_{ij}}$ (4)
where $r_{ij}=M_{j}/M_{i}$. Now for the resonant regime $\Delta M=M_{i}-M_{j}=\Gamma_{0}/2$ then the CP asymmetry becomes maximal $\epsilon\sim\sin{(2\phi)}$ with $\phi$ being the relative phase.
The asymmetry generated in $\nu_{R}$ is then transferred to left-handed leptons via sizeable coupling with $H_{2}$. If this asymmetry is transferred to the left-handed leptons before sphalerons decouple, a non-zero baryon asymmetry can be generated. Neutrinos acquire small Dirac mass at tree level from the vacuum expectation value (VEV) of the neutral component of $H_{2}$ (denoted by $v_{2}$) as
@@
where $v_{2}$ can be much smaller than the VEV of the SM-like Higgs doublet $H_{1}$. This can be achieved if the VEV of $H_{2}$ is protected by an approximate global symmetry and can only be induced after electroweak symmetry breaking due to the VEV of $H_{1}$. Smallness of $v_{2}$ compared to $v_{1}\equiv\langle H_{1}\rangle$ ensures that the Yukawa coupling $y$ is sufficiently large to transfer the asymmetry from $\nu_{R}$ to $\nu_{L}$ as we discuss below.
@@
A large Majorana mass of $\nu_{R}$ can introduce additional source of lepton asymmetry and washout of the asymmetry already produced from $N_{i}$ decay. As this Majorana mass term is very similar to the decay width of $N_{i}$, keeping it small also forces $\Gamma_{i}$ to be small, which leads to interesting correlations among the scale of leptogenesis and light sterile neutrino masses.
@@ II.1 Leptogenesis
The relevant Boltzmann equations can be written as
@@
$K=\frac{\Gamma_{1}(z=\infty)}{\mathcal{H}(z=1)}.$ (10)
@@
The scattering processes leading to washout include $\nu_{R}\nu_{R}\leftrightarrow aa$ as one of the dominant ones. The requirement for keeping this washout process out of equilibrium leads to
@@
$f\gtrsim\sqrt{\Gamma_{1}M_{\rm Pl}}$ (13)
Since $\Lambda>M_{1}$ for the validity of the effective operator, the above condition also ensures $K<1$, keeping the washouts under control. Another important washout process is $LL\leftrightarrow H_{2}H_{2}$ or other variants like $LH^{\dagger}_{2}\leftrightarrow\overline{L}H_{2}$ which violate lepton number by two units. Demanding this washout to be out of equilibrium $(\Gamma^{\rm wo}_{2}<\mathcal{H})$ leads to
@@
which, for $\Gamma_{1}\sim 1$ keV, leads to $y<\mathcal{O}(1)$. For larger decay width of $N_{1}$ and hence larger Majorana mass of $\nu_{R}$, the upper bound will be stronger. The Dirac Yukawa can not be arbitrarily low as the asymmetry in $\nu_{R}$ gets transferred to the left sector via this Yukawa portal interactions only. Demanding this Yukawa interaction to be in equilibrium prior to the sphaleron decoupling epoch leads to a lower bound
@@
which gets stronger if the Yukawa interactions are required to be in equilibrium at higher temperatures. However, as far as the generation of baryon asymmetry is concerned, it is sufficient to ensure that this process enters equilibrium before the sphalerons decouple.
Once the asymmetry in $\nu_{R}$ is transferred to the left-handed lepton doublets, the electroweak sphaleron processes convert the $B-L$ asymmetry into baryon asymmetry as
@@
where $C_{\rm sph}=\frac{8}{23}$ for two Higgs doublets $N_{\textbf{H}}=2$ and three fermion generations $N_{f}=3$. The factor $S$ accounts for the change in the relativistic degrees of freedom from the scale of leptogenesis until recombination and comes out to be $S=\frac{106.75}{3.91}\simeq 27.3$.
@@
Figure 4: Allowed parameter space of minimal ALPy cogenesis scenario in $f-M_{1}$ plane considering dimensionless couplings $\sim 10^{-2}$. The shaded regions are disfavored leaving only the white region allowed for minimal ALPy cogenesis.
Fig. 3 shows the evolution of comoving number densities of heavy RHN, and the absolute value of $\nu_{R}$ asymmetry. As the RHN abundance gradually rises, the lepton asymmetry gets a sign-flip leading to the resonance type feature. The lepton asymmetry saturates subsequently to the required value and transferred into asymmetry in lepton doublets which are in equilibrium with $\nu_{R}$ via neutrinophilic Higgs doublet $H_{2}$. The asymmetry in lepton doublet is then converted into baryon asymmetry when electroweak sphalerons go out of equilibrium at $T=T_{\rm sph}$. Fig. 4 shows the parameter space consistent with successful minimal ALPy cogenesis in $f-M_{1}$ plane for two different values of CP asymmetry parameter $\epsilon_{1}$. The upper left triangular part of the plane labeled as $\Gamma_{\rm agg}/\mathcal{H}<1$ corresponds to the region where axions-gluon interactions are not in thermal equilibrium Masso:2002np . Since we are assuming $a,\nu_{R}$ to be part of the bath while solving the Boltzmann equations, our analysis remains valid outside this shaded region only. The shaded regions labeled as $\Gamma^{\rm wo}_{1,2}/\mathcal{H}>1$ are disfavored due to strong washout rates. The horizontal shaded region at the top correspond to the parameter space where QCD axion DM is overproduced via vacuum misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah .

### v2

@@ I Introduction
The origin of dark matter (DM) and baryon asymmetry in the Universe (BAU) have been longstanding puzzles in particle physics and cosmology Zyla:2020zbs ; Planck:2018vyg . The failure of the standard model (SM) of particle physics to explain these observed phenomena has led to several beyond standard model (BSM) proposals in the literature. Among them, the weakly interacting massive particle (WIMP) paradigm of DM Kolb:1990vq ; Jungman:1995df and baryogenesis/leptogenesis Weinberg:1979bt ; Kolb:1979qa ; Fukugita:1986hr are the most popular ones. While the minimal WIMP scenarios are pushed to a tight corner by null results at direct search experiments LZ:2022lsv , generic baryogenesis models operate at a high scale out of reach from direct experimental probes. Minimal vanilla leptogenesis scenarios also have a lower bound on the scale of leptogenesis $\geq 10^{9}$ GeV, known as the Davidson-Ibarra (DI) bound Davidson:2002qv . While a low scale realization of leptogenesis with resonantly enhanced CP asymmetry Pilaftsis:2003gt exists, some well-motivated DM paradigms beyond the simplest WIMP paradigm have also been explored in the literature. The axion or axion-like particle (ALP) provides such an alternative paradigm for DM, a recent review of which can be found in OHare:2024nmr .
@@
The axion Wilczek:1977pj ; Weinberg:1977ma is the pseudo Nambu-Goldstone boson (pNGB) associated with the Peccei-Quinn (PQ) symmetry $U(1)_{\rm PQ}$ Peccei:1977hh ; Peccei:1977ur ; Wilczek:1977pj ; Weinberg:1977ma introduced to solve the strong CP problem, another longstanding puzzle in particle physics. In addition to solving the strong CP problem, an axion with sub-eV mass can constitute the entire cold dark matter in the Universe Preskill:1982cy ; Abbott:1982af ; Dine:1982ah . Axions can also have interesting detection prospects due to their couplings to the SM particles, suppressed by the PQ symmetry breaking scale $f$. For QCD axion models with sub-eV axion mass, there exists a stringent astrophysical bound $f\gtrsim 10^{8}-10^{9}$ GeV Raffelt:2006cw ; Caputo:2024oqc . On the other hand, ALP parameter space is less constrained as their couplings and mass can be varied independently. With the growing interest in axion DM and its related phenomenology, it is natural to ask whether this DM paradigm alternative to the WIMP has implications for the observed baryon asymmetry of the Universe.
@@
Motivated by these questions, we propose a novel way of connecting leptogenesis and axion by considering a specific interaction involving axion and suppressed by the PQ scale $f$ to be the origin of CP violation. While there are several ways of connecting the axion with baryogenesis or leptogenesis Kusenko:2014uta ; Servant:2014bla ; Adshead:2015jza ; Ipek:2018lhm ; Croon:2019ugf ; Co:2019wyp ; Ballesteros:2016xej ; Clarke:2015bea ; Co:2020jtv ; Sopov:2022bog ; Domcke:2022kfs ; Borah:2024qyo ; Mishra:2025twb ; Foster:2022ajl , we for the first time, propose a decay of the type $\psi_{1}\rightarrow\psi_{2}a$ (with $\psi_{1,2}$ being chiral fermions and $a$ being the axion) to be the origin of the observed CP asymmetry. As a first working example of this mechanism, dubbed as ALPy cogenesis, we consider $\psi_{1}$ to be a heavy right-handed neutrino (RHN) $N_{R}$ while $\psi_{2}$ to be a light chiral fermion $\nu_{R}$, both of which are singlets under the SM gauge symmetry [^1] In another recent work Cataldi:2024bcs , a non-thermal leptogenesis scenario was studied where a heavy ALP decays into RHN in the early Universe. On the other hand, Datta:2024xhg considered spontaneous leptogenesis with sub-GeV ALP.. A non-zero CP asymmetry is generated after taking one-loop correction to this decay if at least two copies of $N_{R}$ are considered. The asymmetry generated in $\nu_{R}$ is then transferred to left-handed leptons via sizeable Yukawa interactions with a neutrinophilic Higgs doublet Heeck:2013vha ; Borah:2022qln . Due to the lepton number violation induced by heavy RHN $N_{R}$, a Majorana mass of $\nu_{R}$ is also generated at one-loop level. This radiative correction splits the three Dirac neutrinos at tree level into three active and three sterile Majorana or pseudo-Dirac neutrinos. Keeping the Majorana mass of $\nu_{R}$ small to control the washout of the asymmetry created from $N_{R}$ decay also constrains the scale of leptogenesis and other parameters of the model. As a second working example, we consider a setup without lepton number violation which ensures light neutrinos to be of Dirac type. This requires the presence of another Dirac fermion $\chi$ such that the net CP asymmetries in $\nu_{R}$ and $\chi$ can be zero due to lepton number conservation. Keeping the $\nu_{R}$ and $\chi$ sectors out of equilibrium allow a net lepton asymmetry to survive till the sphaleron decoupling epoch. The asymmetric $\chi$, if sufficiently long lived, can contribute to DM along with axions. In both the examples, the scale of leptogenesis as well as the PQ scale are constrained from the requirement of generating the observed BAU and DM. The PQ scale, on the other hand, also controls the detection prospects of axion or ALP over a wide range of experiments. The thermal leptogenesis model explored here can lead to multiple sources of dark radiation that, taken together, might be within the reach of current or future cosmic microwave background (CMB) experiments. The dark radiation can include both axions, which may be thermalized via the axion portal operator responsible for generating lepton asymmetry, as well as Dirac neutrinos. It should be noted that the presence of ALP or axion in the interaction of heavy neutral fermion is primarily motivated from enhanced detection prospects of the mechanism. For example, leptogenesis can still work if the axion is replaced by a photon, as studied in the context of electromagnetic leptogenesis Bell:2008fm ; Choudhury:2011gbi ; Borah:2025oqj or by a real scalar.
@@
This paper is organized as follows. In section II, we discuss the minimal framwork with Majorana neutrinos. In section III, we discuss the Dirac ALPy cogenesis scenario followed by the detection prospects of both the scenarios in section IV. We finally conclude in section V.
@@ II Minimal Framework
In the minimal realization of the ALPy cogenesis idea, the SM lepton content is extended by two different types of singlet chiral fermions $N_{R},\nu_{R}$. We consider three copies each for these chiral fermions. While $\nu_{R}$ combines with the left chiral neutrinos $\nu_{L}$ to form Dirac neutrinos at tree level via a neutrinophilic Higgs doublet $H_{2}$, the heavy right-handed neutrino $N_{R}$ has an effective ALP coupling with $\nu_{R}$. The relevant Lagrangian is given by
-$\mathcal{L}$ $\supset\frac{\partial_{\mu}a}{f}\overline{N_{R_{i}}}(\lambda_{i\alpha}\gamma^{\mu}+\lambda^{\prime}_{i\alpha}\gamma^{\mu}\gamma^{5})\nu_{R_{\alpha}}+\frac{1}{2}M_{i}\overline{N^{c}_{R_{i}}}N_{R_{i}}$
@@
Similar higher dimensional operators with axion interaction being replaced by electromagnetic dipole operator were used for leptogenesis in earlier works Bell:2008fm ; Choudhury:2011gbi ; Borah:2025oqj . While we do not study the details of UV complete models in this work, suitable discrete symmetries can be imposed to realize the above interactions. For example, a $Z_{4}$ symmetry with the above-mentioned fields charged as $N_{R_{i}}(-1),\nu_{R}(i),H_{2}(i),a(i)$ can prevent renormalizable Yukawa interactions of $N_{R_{i}}$ with the SM leptons either via SM-like Higgs $H_{1}$ or neutrinophilic Higgs $H_{2}$. Additionally, a soft $Z_{4}$ breaking term in the scalar potential $\mu^{2}_{\rm soft}H^{\dagger}_{1}H_{2}$ leads to an induced vacuum expectation value (VEV) of $H_{2}$ as
$\langle H^{0}_{2}\rangle\equiv v_{2}\approx\frac{\mu^{2}_{\rm soft}v_{1}}{M^{2}_{H_{2}}}$ (2)
where $v_{1}$ is the VEV of the SM-like Higgs $H_{1}$ and $M_{H_{2}}$ is the bare mass of $H_{2}$. This also leads to mixing between neutral scalar of $H_{2}$ and the SM Higgs resulting in interesting phenomenology which has been studied in several earlier works Maitra:2014qea ; Haba:2011nb ; Gabriel:2008es ; Seto:2015rma ; Bertuzzo:2015ada ; Huitu:2017vye .
Fig. 1 shows the two-body decay $N_{i}\rightarrow\nu_{R_{\alpha}}a$ and possible one-loop corrections responsible for generating the CP asymmetry. The corresponding CP asymmetry, where the lepton flavor $\alpha$ of $\nu_{R_{\alpha}}$ is being summed over, can be written as
@@
$\Lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz,\\$
and $r_{a}=m_{a}/M_{i},r_{\alpha,\beta}=m_{\nu_{\alpha,\beta}}/M_{i}$, $\widetilde{\lambda}_{i\alpha}=\lambda_{i\alpha}+\lambda^{\prime}_{i\alpha}$. We always work in the limit where $r_{a},r_{\alpha}\ll 1$ such that $\Gamma_{0}\approx\frac{|\widetilde{\lambda}_{i\alpha}|^{2}M_{i}}{32\pi}\left(\frac{M_{i}}{f}\right)^{2}$. The CP asymmetry parameter, considering self-energy correction, is given by
$\epsilon_{i}=\frac{{\rm Im}[\widetilde{\lambda}_{j\alpha}\widetilde{\lambda}_{j\beta}\widetilde{\lambda}^{*}_{i\beta}\widetilde{\lambda}^{*}_{i\alpha}]}{|\widetilde{\lambda}_{i\beta}|^{2}|\widetilde{\lambda}_{i\alpha}|^{2}M_{i}}\Gamma_{0}(N_{i}\rightarrow\nu_{R\beta}a)\frac{r_{ij}}{1-r^{2}_{ij}}$ (5)
where $r_{ij}=M_{j}/M_{i}$. Now for the resonant regime $\Delta M=M_{i}-M_{j}=\Gamma_{0}/2$ then the CP asymmetry becomes maximal $\epsilon\sim\sin{(2\phi)}$ with $\phi$ being the relative phase. In order to arrive at this expression, $\lvert\tilde{\lambda}\rvert$ is assumed to be the common magnitude of the couplings $\tilde{\lambda}_{i\alpha}=\lvert\tilde{\lambda}\rvert e^{i\theta_{i\alpha}}$ with a common relative phase angle given by $\phi=\theta_{i\alpha}-\theta_{j\alpha}$.
The asymmetry generated in $\nu_{R}$ is then transferred to the left-handed leptons via sizeable coupling with $H_{2}$. If this asymmetry is transferred to the left-handed leptons before sphalerons decouple, a non-zero baryon asymmetry can be generated. The neutrinos acquire small Dirac mass at tree level from the VEV of the neutral component of $H_{2}$ as
@@
where $v_{2}$ can be much smaller than the VEV of the SM-like Higgs doublet $H_{1}$. This can be achieved if the VEV of $H_{2}$ is protected by an approximate global symmetry like the $Z_{4}$ symmetry mentioned above and can only be induced after electroweak symmetry breaking due to the VEV of $H_{1}$. Smallness of $v_{2}$ compared to $v_{1}\equiv\langle H_{1}\rangle$ ensures that the Yukawa coupling $y$ is sufficiently large to transfer the asymmetry from $\nu_{R}$ to $\nu_{L}$ as we discuss below.
@@
A large Majorana mass of $\nu_{R}$ can introduce additional source of lepton asymmetry and washout of the asymmetry already produced from $N_{i}$ decay. As this Majorana mass term is very similar to the decay width of $N_{i}$, keeping it small also forces $\Gamma_{i}$ to be small, which leads to interesting correlations among the scale of leptogenesis and light sterile neutrino masses. Note that this similarity between decay width of $N_{i}$ and Majorana mass of $\nu_{R}$ holds only in the absence of any specific texture of $\tilde{\lambda}$ or strong phase-space suppression in $\Gamma_{0}$. In our analysis, we are neither using any specific textures nor phase-space suppression and we explicitly calculate them without assuming any similarities.
@@ II.1 Leptogenesis
While there are three heavy neutral fermions $N_{i}$, the lepton asymmetry is generated by out-of-equilibrium decay of the lightest among them. The relevant Boltzmann equations can be written as
@@
$K=\frac{\Gamma_{1}}{\mathcal{H}(z=1)}.$ (11)
@@
The scattering processes leading to washout include $\nu_{R}\nu_{R}\leftrightarrow aa$ as one of the dominant ones. The requirement for keeping this washout process (mediated by $N_{1}$) out of equilibrium leads to
@@
$f\gtrsim\sqrt{\Gamma_{1}M_{\rm Pl}}.$ (14)
For $f>M_{1}$, the above condition also ensures $K<1$, keeping the washout under control. Another important washout process is $LL\leftrightarrow H_{2}H_{2}$ or other variants like $LH^{\dagger}_{2}\leftrightarrow\overline{L}H_{2}$ which violate lepton number by two units. Demanding this washout to be out of equilibrium $(\Gamma^{\rm wo}_{2}<\mathcal{H})$ leads to
@@
which, for $\Gamma_{1}\sim 1$ keV, leads to $y<\mathcal{O}(1)$. For larger decay width of $N_{1}$ and hence larger Majorana mass of $\nu_{R}$, the upper bound will be stronger. The Dirac Yukawa coupling can not be arbitrarily small as the asymmetry in $\nu_{R}$ gets transferred to the left sector via this Yukawa portal interactions only. Demanding this Yukawa interaction to be in equilibrium prior to the sphaleron decoupling epoch leads to a lower bound
@@
which gets stronger if the Yukawa interactions are required to be in equilibrium at higher temperatures. However, as far as the generation of baryon asymmetry is concerned, it is sufficient to ensure that this process enters equilibrium before the sphalerons decouple. We are also not assuming any specific flavor structure of the Dirac Yukawa coupling $y$ in our analysis, assuming it to be flavor universal. A full treatment of flavored leptogenesis Abada:2006fw ; Nardi:2006fx ; Abada:2006ea ; Blanchet:2006be ; BhupalDev:2014pfm ; Dev:2017trv is beyond the scope of the present work.
Once the asymmetry in $\nu_{R}$ is transferred to the left-handed lepton doublets, the electroweak sphaleron processes convert the BLB-L asymmetry into baryon asymmetry as
@@
where $C_{\rm sph}=\frac{8}{23}$ for two Higgs doublets $N_{\textbf{H}}=2$ and three fermion generations $N_{f}=3$. The factor $S$ accounts for the change in the relativistic degrees of freedom from the scale of leptogenesis until recombination and comes out to be $S=\frac{106.75}{3.91}\simeq 27.3$. The final baryon asymmetry $\eta_{\rm B}$ can also be estimated analytically in terms of the CP asymmetry parameter as Buchmuller:2004nz
$\eta_{B}=\frac{C_{\rm sph}}{S}\epsilon_{1}\kappa\,$, (19)
where $\kappa$ is known as the efficiency factor which incorporates the effects of washout processes. Instead of considering any approximate value of $\kappa$ valid for a specific washout regime, we solve the Boltzmann equations numerically to calculate the asymmetry explicitly.
@@
Figure 4: Allowed parameter space of minimal ALPy cogenesis scenario in $f-M_{1}$ plane considering dimensionless couplings $\tilde{\lambda}_{i\alpha},y_{\alpha\beta}\sim 10^{-2}$. The shaded regions are disfavored leaving only the white region allowed for minimal ALPy cogenesis.
Fig. 3 shows the evolution of comoving number densities of heavy RHN, and the absolute value of $\nu_{R}$ asymmetry. As the RHN abundance gradually rises, the lepton asymmetry gets a sign-flip leading to the resonance type feature. The lepton asymmetry saturates subsequently to the required value and transferred into asymmetry in lepton doublets which are in equilibrium with $\nu_{R}$ via neutrinophilic Higgs doublet $H_{2}$. The asymmetry in lepton doublet is then converted into baryon asymmetry when electroweak sphalerons go out of equilibrium at $T=T_{\rm sph}$. Fig. 4 shows the parameter space consistent with successful minimal ALPy cogenesis in $f-M_{1}$ plane for two different values of the CP asymmetry parameter $\epsilon_{1}$. The points on the blue and maroon colored solid lines are consistent with the observed baryon asymmetry for two different values of the CP asymmetry parameter $\epsilon_{1}=3.67\times 10^{-7},\epsilon_{1}=10^{-2}$ respectively. The upper left triangular part of the plane labeled as $\Gamma_{\rm agg}/\mathcal{H}<1$ corresponds to the region where the axion-gluon interactions are not in thermal equilibrium Masso:2002np . Since we are assuming $a,\nu_{R}$ to be part of the bath while solving the Boltzmann equations, our analysis remains valid outside this shaded region only. The calculation of the baryon asymmetry with non-thermal axions require solving the Boltzmann equation for axions and estimating the corresponding non-thermal abundance of heavy neutral fermions. We do not pursue such an analysis here and stick to the thermal leptogenesis regime only. The shaded regions labeled as $\Gamma^{\rm wo}_{1,2}/\mathcal{H}>1$ are disfavored due to strong washout rates. The horizontal shaded region at the top corresponds to the parameter space where the QCD axion DM is overproduced via the vacuum misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah . It should be noted that the dimensionless couplings $\tilde{\lambda}_{i\alpha},y_{\alpha\beta}$ are assumed to be $\sim 10^{-2}$ in Fig. 4. The constraints from washouts become stronger for larger values of these dimensionless couplings, shrinking the allowed white colored region in Fig. 4.